use an algebraic equation to solve:
the sides of a triangle are in the ratio 3:4:5. what is the length of each side if the perimeter is 30cm?

Answer by ankor@dixie-net.com(22740)

he sides of a triangle are in the ratio 3:4:5.
what is the length of each side if the perimeter is 30cm?
:
let x = the multiplier
then
3x + 4x + 5x = 30
12x = 30
x = 30/12
x = 2.5 if the multiplier, therefore the three sides are:
3*2.5 = 7.5
4*2.5 = 10
5*2.5 = 12.5